# Excitation Spectrum and Spin Hamiltonian of the Frustrated Quantum Ising Magnet PrBWO

### v1

@@ III.9 Crystal Structure & Disorder
Given the success of our disorder model in explaining both the thermodynamics and spectrum of PrBWO, it is well worth delving into the potential causes for disorder by taking a closer look at the crystal structure. To that end, we collect neutron powder diffraction data on B-substituted samples and carry out Rietveld refinements, the results of which are shown in Figure 12(a). The final -factor including anisotropic thermal parameters amounts to and the refined atomic positions (utilized for the point-charge calculations in Sec. III.3) are listed in Table 6.
@@
Perhaps the most common source of structural disorder found in quantum magnets is antisite mixing, a random switching of ions with similar atomic radii which is notorious for its role in spin-liquid candidates such as Herbertsmithite [hanCorrelatedImpuritiesIntrinsic2016] or YbMgGaO [liCrystallineElectricFieldRandomness2017, zhuDisorderInducedMimicrySpin2017]. The large differences between atomic radii makes this a rather unlikely problem in PrBWO [ashtarNewFamilyDisorderFree2020], which is confirmed by our refinement attempts, showing no signs of antisite disorder within uncertainties (%).
@@
Another common concern is off-stoichiometry, as seen e.g. in pyrochlore oxides [maclaughlinUnstableSpiniceOrder2015, koohpayehSynthesisFloatingZone2014]. A refinement with variable site occupancy indeed points to a slight excess in oxygen concentration of compared to the other ions, yielding an improvement in -factors. In Pr-based oxides, the charge imbalance created by the extra O is typically accommodated by increasing the valence for some fraction of magnetic ions from Pr to Pr [koohpayehSynthesisFloatingZone2014, ferroPhysicochemicalElectricalProperties2011]. Even though all powders were sintered for several days at 1200C during synthesis until a uniform green color was obtained (indicative of pure Pr valence [ferroPhysicochemicalElectricalProperties2011]) and the XRD spectra no longer changed, it would seem that some trace amounts of Pr and O impurities have remained throughout this process.
@@
If our samples really are contaminated with magnetic Pr impurities, these should give rise to an observable Curie-like tail in the low temperature susceptibility. To test this hypothesis, we compare magnetic susceptibility curves collected on several different single crystals. Subtracting from two samples grown from different batches as depicted in Figure 12(b,c), the difference can indeed be fit to a contribution, amounting roughly to a concentration of impurity spins. Although this number corresponds to the difference between Pr concentrations from two separate samples, we can also establish an upper bound on the absolute number of impurity spins. This is because the Van-Vleck susceptibility of a singlet ground state system for is constant - i.e. the susceptibility should not have a positive slope after subtracting a potential Curie-like impurity tail. A significantly smaller concentration than that obtained from powder diffraction is to be expected, as single crystal samples are commonly of higher quality than powders.
@@
With this, we have confirmation from two separate sources - neutron powder diffraction and magnetic susceptibility measurements - that a small surplus of oxygen and the concomitant Pr charge impurities are responsible for the structural disorder detected in other probes. Although might seem like a small amount, the local distortions resulting from such disorder decay as a power-law with distance. So even though relatively few magnetic ions are directly adjacent to an impurity site, the vast majority of sites will be “affected” by the long-ranged structural distortions emanating from the local impurity sites, creating small modifications to the local crystal electric field. And while such corrections would be irrelevant for most common magnets, the lack of Kramers degeneracy makes PrBWO particularly susceptible to this effect. As such, the seemingly huge fraction of sites contributing to the broad continuum scattering discussed in Sec. III.8 can easily be explained by a small number of impurities .
@@ IV Discussion
While PrBWO was initially advertised as a proximate spin-liquid candidate on the breathing kagome lattice [ashtarNewFamilyDisorderFree2020, zengLocalEvidenceCollective2021], a more careful investigation reveals a completely disparate picture. We combine thermodynamic, magnetometric, neutron- and muon spectroscopy data with numerical modeling, providing strong evidence that PrBWO can be understood as a singlet ground state system with a small gap meV separating the lowest two crystal field levels. After a pseudospin projection down to this lowest quasi-doublet, the effective Hamiltonian takes the form of a nearly ideal Ising model in a transverse magnetic field. Despite the trivial point symmetry and complicated structure, the effects of crystal field and non-Kramers moments with large- conspire to form a simple model Hamiltonian of significant interest.
@@
The phase diagram was thoroughly characterized, proving that PrBWO remains disordered in the entire phase space. Although both the triangle-based lattice topology and the FM-AFM exchange competition contribute to a significant frustration, neither is the main cause of the quantum disordered ground state. Even on a mean field level with no frustration, the exchange interactions barely amount to of the critical ratio
@@
necessary for induced-moment order [thalmeierInducedQuantumMagnetism2023, jensenRareEarthMagnetism1991], where refers to the Fourier transformed exchange matrix - driving home the point that PrBWO is really a quantum paramagnet and not a quantum spin liquid candidate. In the TFIM pseudospin language, we are situated in the limit of large transverse fields, meaning that the ground state can be understood as an induced ferromagnet, fully polarized by the transverse field produced by the crystal field splitting.
@@
Given these circumstances, linear spin-wave theory constitutes an excellent approximation and can be used to extract the “real” (non-renormalized) exchange parameters of PrBWO with high fidelity [coldeaDirectMeasurementSpin2002]. But while LSWT fits to our neutron spectra establish that significant correlations in the purported kagome planes are present, they are clearly dwarfed by the dominant out-of-plane interactions, forming frustrated spin-tubes through two competing FM-AFM bonds. Although a sign change for (counter-)clockwise nearest-neighbor bonds might seem surprising at first glance given the visually similar Pr-O-Pr superexchange pathways, it is easy to rationalize as stemming from a difference in bond angles, which are rather close to the FM-AFM crossover expected from Kanamori-Goodenough rules [goodenoughMagnetismChemicalBond]. Furthermore, it is interesting to note here the relative differences between planar/non-planar exchange constants in the full model and the effective Hamiltonian: While the former would indicate a 3D exchange topology with , the latter points to a nearly one-dimensional physics with . This apparent discrepancy is fully accounted for by the pseudospin projection. The easy axes between the Ising-like moments forming our breathing kagome triangles are nearly orthogonal, leading to an effective suppression proportional to of the in-plane couplings. While an accurate characterization of the breathing anisotropy is complicated by this suppression, we estimate based on our LSWT calculations - not surprising given the significant () difference in bond length and the highly localized character of orbitals.
@@
But despite the obvious contrast between the supposed, highly degenerate breathing kagome model and the simpler, nearly 1D two-singlet physics at hand, PrBWO exhibits a number of puzzling properties that can only be understood by considering two important extensions to the pseudospin-projected TFIM Hamiltonian.
@@
Furthermore, due to the non-Kramers nature of the Pr ion, the material is especially susceptible to effects of structural disorder, which results in a distribution of CEF splittings . Even though our single crystal samples display only a meager impurity concentration (in the form of Pr charge disorder and a concomitant excess of oxygen), there are dramatic effects on both the thermodynamic and dynamic properties. The heat capacity becomes with a power-law decay towards low temperatures, while the bulk of the dispersive crystal field excitons are restructured into a broad of excitations. We note that while both of these features are commonly provided as evidence for spin-liquid like physics, the extreme sensitivity of the magnetism to even small amounts of disorder seen here reinforces our belief that effects of structural imperfections should be carefully considered before claiming such exotic physics, especially when dealing with non-Kramers systems. Also, the coexistence of both sharp spin-waves and a broad continuum is particularly unusual in disordered magnets - most of the known material examples with considerable structural disorder exhibit a broadened spectrum [hanCorrelatedImpuritiesIntrinsic2016, liCrystallineElectricFieldRandomness2017, liSpinDynamicsGriffiths2021], not two separate components. While it is possible to interpret the sharp component in terms of local regions sufficiently removed/screened from impurities, a significant variation of impurity concentrations between the coaligned single crystals seems to be a much more likely explanation, especially considering the variation already found in susceptibility measurements.
@@
There is one more subtle feature which might originate from structural disorder, namely the rapid temperature-dependent softening of the sound velocity at low fields T. Any magnetoelastic effects associated with the Pr impurity spins should be suppressed at modest fields , potentially explaining the otherwise anomalously small energy scale associated with this feature.
@@
Finally, we compare PrBWO to other members in the rare-earth borotungstate family. The physics of Ising-like moments coupled into 1D frustrated spin-tubes arises primarily from the local environment, so given the strong structural similarities, we would expect this framework to remain intact for most of the light rare earths, as evidenced e.g. in [flavianMagneticPhaseDiagram2023a]. Obvious exceptions to this would be (a sign change of the Stevens factor should provide a strong easy plane anisotropy [jensenRareEarthMagnetism1991]), as well as Ce (the smaller allows for significant -type exchange components despite the Ising anisotropy [iwaharaExchangeInteractionMultiplets2015]). While the effects of both planar terms and unsuppressed in-plane exchanges might be interesting to study, the two-ion Hamiltonian should be decidedly more complicated due to the low symmetry. Also, judging from the twofold enhanced ordering temperature of Sm (planar) [zengIncommensurateMagneticOrder2022] compared to Nd (axial) [flavianMagneticPhaseDiagram2023a], it would seem that the frustration might be partially released by the inverted anisotropy.
@@
In the context of the TFIM physics seen here, an exploration of the members might be of particular interest. While PrBWO unfortunately ends up in a pseudospin-polarized region of phase space with a transverse field exceeding the exchanges, the crystal field scale for the latter ions should be significantly smaller [jensenRareEarthMagnetism1991]. This opens up the question how the interplay between geometric/exchange frustration and transverse field-induced quantum fluctuations might modify the ground state for smaller fields - especially whether a strongly correlated quantum disordered phase can remain stable in an extended regime . Potential structural disorder should also be more subdued in these materials, though its effects might even be beneficial, having been suggested to stabilize a quantum spin liquid phase in certain frustrated systems [savaryDisorderInducedQuantumSpin2017].
@@ V Conclusion
Despite the promise of spin-liquid physics on a breathing kagome lattice, PrBWO turns out to be a relatively simple quantum paramagnet. While it does feature significant frustration inherent from both the lattice topology and an FM-AFM exchange competition, the ground state properties are dictated by a crystal field splitting of the lowest quasi-doublet. This results in a simple effective Hamiltonian of a frustrated Ising model in a transverse magnetic field, which can be modeled numerically to astounding success. Our neutron spectroscopy results demonstrate that the dominant interactions are those perpendicular to the kagome planes, confirming that the BWO family of quantum antiferromagnets is best described as an arrangement of 1D frustrated triangular spin-tubes, moderately coupled through the planar breathing kagome exchanges. The two-singlet nature of PrBWO makes it especially susceptible to the influence of hyperfine enhanced nuclear moments and weak structural disorder, both of which have dramatic effects on the static and dynamic properties, including an unusual coexistence of both sharp spin-waves and broad continuum excitations in an otherwise simple model system. Frustrated quantum Ising models remain a promising avenue in the ongoing hunt for novel emergent quantum magnetic phases, although the desired parametric regime of frustrated exchanges and strong quantum fluctuations continues to prove rather difficult to engineer.
@@ VI Acknowledgements
Appendix A: Effective Point-Charge Model
( meV) |
Global Frame |
Local Frame |
-45.732 |
-24.513 |
-464.458 |
-439.061 |
30.383 |
45.214 |
-358.590 |
-288.734 |
-86.931 |
-262.987 |
-0.776 |
-3.805 |
-11.341 |
6.819 |
-0.786 |
0.777 |
6.168 |
11.124 |
1.121 |
1.708 |
-7.957 |
-5.627 |
-1.515 |
-2.545 |
-1.005 |
-15.699 |
2.920 |
-3.648 |
-0.114 |
0.119 |
0.099 |
0.264 |
0.068 |
0.120 |
0.020 |
-0.044 |
-0.057 |
-0.066 |
0.066 |
0.096 |
0.002 |
0.006 |
-0.060 |
-0.007 |
0.023 |
-0.044 |
0.008 |
0.051 |
-0.132 |
0.172 |
-0.444 |
0.451 |
-0.024 |
0.025 |

### v2

@@ III.9 Crystal Structure & Disorder
Given the success of our disorder model in explaining both the thermodynamics and spectrum of Pr3BWO9, it is well worth delving into the potential causes for disorder by taking a closer look at the crystal structure. To that end, we collect neutron powder diffraction data on 11B-substituted samples and carry out Rietveld refinements, the results of which are shown in Figure 12(a). The final -factor including anisotropic thermal parameters amounts to and the refined atomic positions (utilized for the point-charge calculations in Sec. III.3) are listed in Table 6.
@@
Perhaps the most common source of structural disorder found in quantum magnets is antisite mixing, a random switching of ions with similar atomic radii which is notorious for its role in spin-liquid candidates such as Herbertsmithite [hanCorrelatedImpuritiesIntrinsic2016] or YbMgGaO4 [liCrystallineElectricFieldRandomness2017, zhuDisorderInducedMimicrySpin2017]. The large differences between atomic radii makes this a rather unlikely problem in Pr3BWO9 [ashtarNewFamilyDisorderFree2020], which is confirmed by our refinement attempts, showing no signs of antisite disorder within uncertainties (%).
@@
Another common concern is off-stoichiometry, as seen e.g. in pyrochlore oxides [maclaughlinUnstableSpiniceOrder2015, koohpayehSynthesisFloatingZone2014]. A refinement with variable site occupancy indeed points to a slight excess in oxygen concentration of compared to the other ions, yielding an improvement in -factors. In Pr-based oxides, the charge imbalance created by the extra O2- is typically accommodated by increasing the valence for some fraction of magnetic ions from Pr3+ to Pr4+ [koohpayehSynthesisFloatingZone2014, ferroPhysicochemicalElectricalProperties2011]. Even though all powders were sintered for several days at 1200∘C during synthesis until a uniform green color was obtained (indicative of pure Pr3+ valence [ferroPhysicochemicalElectricalProperties2011]) and the XRD spectra no longer changed, it would seem that some trace amounts of Pr4+ and O2- impurities have remained throughout this process.
@@
If our samples really are contaminated with magnetic Pr4+ impurities, these should give rise to an observable Curie-like tail in the low temperature susceptibility. To test this hypothesis, we compare magnetic susceptibility curves collected on several different single crystals. Subtracting from two samples grown from different batches as depicted in Figure 12(b,c), the difference can indeed be fit to a contribution, amounting roughly to a concentration of impurity spins. Although this number corresponds to the difference between Pr4+ concentrations from two separate samples, we can also establish an upper bound on the absolute number of impurity spins. This is because the Van-Vleck susceptibility of a singlet ground state system for is constant - i.e. the susceptibility should not have a positive slope after subtracting a potential Curie-like impurity tail. A significantly smaller concentration than that obtained from powder diffraction is to be expected, as single crystal samples are commonly of higher quality than powders.
@@
With this, we have confirmation from two separate sources - neutron powder diffraction and magnetic susceptibility measurements - that a small surplus of oxygen and the concomitant Pr4+ charge impurities are responsible for the structural disorder detected through other probes. In elasticity theory, the strain fields produced by such point defects result in lattice distortions that decay as a power law with distance [TeodosiuElasticPointDefect1982] – analogous to the underlying Coulomb forces at the root of these displacements. While the resulting change in CEF parameters is difficult to estimate – requiring detailed knowledge on the atomic displacements – impurities should clearly have long-ranged effects. So although might seem like a small amount of disorder, the vast majority of sites will be “affected” by the long-ranged structural distortions emanating from the local impurity sites, creating small modifications to the local crystal electric field. Such corrections would be irrelevant for most common magnets, but the lack of Kramers degeneracy makes Pr3BWO9 particularly susceptible to this effect. As such, the seemingly huge fraction of sites contributing to the broad continuum scattering discussed in Sec. III.8 can be naturally explained by a small number of impurities .
@@ IV Discussion
While Pr3BWO9 was initially advertised as a proximate spin-liquid candidate on the breathing kagome lattice [ashtarNewFamilyDisorderFree2020, zengLocalEvidenceCollective2021], a more careful investigation reveals a completely disparate picture. We combine thermodynamic, magnetometric, neutron- and muon spectroscopy data with numerical modeling, providing strong evidence that Pr3BWO9 can be understood as a singlet ground state system with a small gap meV separating the lowest two crystal field levels. After a pseudospin projection down to this lowest quasi-doublet, the effective Hamiltonian takes the form of a nearly ideal Ising model in a transverse magnetic field. Despite the trivial point symmetry and complicated structure, the effects of crystal field and non-Kramers moments with large- conspire to form a simple model Hamiltonian of significant interest.
@@
The phase diagram was thoroughly characterized, proving that Pr3BWO9 remains disordered in the entire phase space. Although both the triangle-based lattice topology and the FM-AFM exchange competition contribute to a significant magnetic frustration, neither is the main cause of the quantum disordered ground state. Even if we completely disregard frustration effects, the exchange interactions barely amount to of the critical ratio
@@
necessary for induced-moment order on a mean-field level [thalmeierInducedQuantumMagnetism2023, jensenRareEarthMagnetism1991], where refers to the Fourier transformed exchange matrix - driving home the point that Pr3BWO9 is really a quantum paramagnet and not a quantum spin liquid candidate. In the TFIM pseudospin language, we are situated in the limit of large transverse fields, meaning that the ground state can be understood as an induced ferromagnet, fully polarized by the transverse field produced by the crystal field splitting.
@@
Given these circumstances, linear spin-wave theory constitutes an excellent approximation and can be used to extract the “real” (non-renormalized) exchange parameters of Pr3BWO9 with high fidelity [coldeaDirectMeasurementSpin2002]. But while LSWT fits to our neutron spectra establish that significant correlations in the purported kagome planes are present, they are clearly dwarfed by the dominant out-of-plane interactions, forming frustrated spin-tubes through two competing FM-AFM bonds. Although a sign change for (counter-)clockwise nearest-neighbor bonds might seem surprising at first glance given the visually similar Pr-O-Pr superexchange pathways, it is easy to rationalize as stemming from a difference in bond angles, which are rather close to the FM-AFM crossover expected from Kanamori-Goodenough rules [goodenoughMagnetismChemicalBond]. Furthermore, it is interesting to note here the relative differences between planar/non-planar exchange constants in the full CEF model and the effective TFIM Hamiltonian: While the former would indicate a 3D exchange topology with , the latter points to a nearly one-dimensional physics with . This apparent discrepancy is fully accounted for by the pseudospin projection. The easy axes between the Ising-like moments forming our breathing kagome triangles are nearly orthogonal, leading to an effective suppression proportional to of the in-plane couplings. While an accurate characterization of the breathing anisotropy is complicated by this suppression, we estimate based on our LSWT calculations - not surprising given the significant () difference in bond length and the highly localized character of orbitals.
@@
But despite the obvious contrast between the supposed, highly degenerate breathing kagome model and the simpler, nearly 1D two-singlet physics at hand, Pr3BWO9 exhibits a number of puzzling properties that can only be understood by considering two important extensions to the pseudospin-projected TFIM Hamiltonian.
@@
Furthermore, due to the non-Kramers nature of the Pr3+ ion, the material is especially susceptible to effects of structural disorder, which results in a distribution of CEF splittings . Even though our single crystal samples display only a meager impurity concentration (in the form of Pr4+ charge disorder and a concomitant excess of oxygen), there are dramatic effects on both the thermodynamic and dynamic properties. The heat capacity becomes with a power-law decay towards low temperatures, while the bulk of the dispersive crystal field excitons are restructured into a broad of excitations. We note that while both of these features are commonly provided as evidence for spin-liquid like physics, the extreme sensitivity of the magnetism to even small amounts of disorder seen here reinforces our belief that effects of structural imperfections should be carefully considered before claiming such exotic physics, especially when dealing with non-Kramers systems. Also, the coexistence of both sharp spin-waves and a broad continuum is particularly unusual in disordered magnets - most of the known material examples with considerable structural disorder exhibit a broadened spectrum [hanCorrelatedImpuritiesIntrinsic2016, liCrystallineElectricFieldRandomness2017, liSpinDynamicsGriffiths2021], not two separate components. While it is possible to interpret the sharp component in terms of local regions sufficiently removed/screened from impurities, a significant variation of impurity concentrations between the coaligned single crystals seems to be a much more likely explanation, especially considering the variation already found in susceptibility measurements.
@@
There is one more subtle feature which might originate from structural disorder, namely the rapid temperature-dependent softening of the sound velocity at low fields T. Any magnetoelastic effects associated with the Pr4+ impurity spins should be suppressed at modest fields , potentially explaining the otherwise anomalously small energy scale associated with this feature.
@@
Finally, we compare Pr3BWO9 to other members in the rare-earth borotungstate family. The physics of Ising-like moments coupled into 1D frustrated spin-tubes arises primarily from the local environment, so given the strong structural similarities, we would expect this framework to remain intact for most of the light rare earths, as evidenced e.g. in [flavianMagneticPhaseDiagram2023a]. Obvious exceptions to this would be (a sign change of the Stevens factor should provide a strong easy plane anisotropy [jensenRareEarthMagnetism1991]), as well as Ce (the smaller allows for significant -type exchange components despite the Ising anisotropy [iwaharaExchangeInteractionMultiplets2015]). While the effects of both planar terms and unsuppressed in-plane exchanges might be interesting to study, the two-ion Hamiltonian should be decidedly more complicated due to the low symmetry. Also, judging from the twofold enhanced ordering temperature of Sm (planar) [zengIncommensurateMagneticOrder2022] compared to Nd (axial) [flavianMagneticPhaseDiagram2023a], it would seem that the frustration might be partially released by the inverted anisotropy.
@@
In the context of the TFIM physics seen here, an exploration of the members might be of particular interest. While Pr3BWO9 unfortunately ends up in a pseudospin-polarized region of phase space with a transverse field exceeding the exchanges, the crystal field scale for the latter ions should be significantly smaller [jensenRareEarthMagnetism1991]. This opens up the question how the interplay between geometric/exchange frustration and transverse field-induced quantum fluctuations might modify the ground state for smaller fields - especially whether a strongly correlated quantum disordered phase can remain stable in an extended regime . Potential structural disorder should also be more subdued in these materials, though its effects might even be beneficial, having been suggested to stabilize a quantum spin liquid phase in certain frustrated systems [savaryDisorderInducedQuantumSpin2017].
@@ V Conclusion
Despite the promise of spin-liquid physics on a breathing kagome lattice, Pr3BWO9 turns out to be a relatively simple quantum paramagnet. While it does feature significant frustration inherent from both the lattice topology and an FM-AFM exchange competition, the ground state properties are dictated by a crystal field splitting of the lowest quasi-doublet. This results in a simple effective Hamiltonian of a frustrated Ising model in a transverse magnetic field, which can be modeled numerically to astounding success. Our neutron spectroscopy results demonstrate that the dominant interactions are those perpendicular to the kagome planes, confirming that the BWO9 family of quantum antiferromagnets is best described as an arrangement of 1D frustrated triangular spin-tubes, moderately coupled through the planar breathing kagome exchanges. The two-singlet nature of Pr3BWO9 makes it especially susceptible to the influence of hyperfine enhanced nuclear moments and weak structural disorder, both of which have dramatic effects on the static and dynamic properties, including an unusual coexistence of both sharp spin-waves and broad continuum excitations in an otherwise simple model system. Frustrated quantum Ising models remain a promising avenue in the ongoing hunt for novel emergent quantum magnetic phases, although the desired parametric regime of frustrated exchanges and strong quantum fluctuations continues to prove rather difficult to engineer.
@@ VI Acknowledgements
Appendix A Effective Point-Charge Model
Following the method pioneered by Hutchings [hutchingsPointChargeCalculationsEnergy1964], we can approximate the crystal field Hamiltonian Eq. (1) by treating the ligands surrounding our magnetic ions as point charges and simply calculating the electrostatic potential. The CEF parameters can be expressed as
where both and are tabulated factors of the tesseral harmonics, is the expectation value of the radial wavefunction [EdvardsonElectrostaticCEF1998] and represent the Stevens factors. As such, we can express Eq. (1) purely in terms of the charges corresponding to the three inequivalent O2- ligands, dramatically reducing the number of free parameters.
| ( meV) | Global Frame | Local Frame |
| -45.732 | -24.513 | |
| -464.458 | -439.061 | |
| 30.383 | 45.214 | |
| -358.590 | -288.734 | |
| -86.931 | -262.987 | |
| -0.776 | -3.805 | |
| -11.341 | 6.819 | |
| -0.786 | 0.777 | |
| 6.168 | 11.124 | |
| 1.121 | 1.708 | |
| -7.957 | -5.627 | |
| -1.515 | -2.545 | |
| -1.005 | -15.699 | |
| 2.920 | -3.648 | |
| -0.114 | 0.119 | |
| 0.099 | 0.264 | |
| 0.068 | 0.120 | |
| 0.020 | -0.044 | |
| -0.057 | -0.066 | |
| 0.066 | 0.096 | |
| 0.002 | 0.006 | |
| -0.060 | -0.007 | |
| 0.023 | -0.044 | |
| 0.008 | 0.051 | |
| -0.132 | 0.172 | |
| -0.444 | 0.451 | |
| -0.024 | 0.025 |
By diagonalizing this CEF Hamiltonian, we obtain the energies and wavefunctions of a single Pr3+ ion, which may be used to fit and compare to experimental data, namely INS and magnetometry. The powder averaged neutron scattering intensity in the dipole approximation is given as
where is a scale factor, represents the magnetic form factor, is the Debye-Waller factor and corresponds to the Boltzmann weights. In practice, the delta function is replaced by a Voigt profile to account for the instrumental resolution and the finite lifetime of excitations, while an intensity correction is introduced to deal with the energy transfer-dependent resolution volume of the spectrometer [PopoviciResolution1975]. On the other hand, the single-ion magnetization is calculated as
where the uniform susceptibility can be simply evaluated numerically as
To recover the bulk magnetic response observed in experiment, these quantities are averaged over all six ions in the unit cell (related by sixfold rotations in the hexagonal plane), i.e.
In the low temperature regime where two-ion correlations start to become relevant, corrections to Eq. (15-17) can be mostly accounted for by a Weiss molecular field. Each ion experiences a local field proportional to the magnetization of its neighbors, determined self-consistently using
Here, is the unit cell volume and corresponds to the unit cell-averaged -tensor, leaving only the mean-field exchange constant as a tuning parameter to deal with interactions.